OpenStudy (anonymous):

How do you simplify: 4a-2/2a^2-a The answer must be 2/a but I don't know how to get to the answer

OpenStudy (anonymous):

yur question not clear ?

OpenStudy (anonymous):

If it's supposed to be \[ \frac{4a - 2}{2a^2 -a} \] Then \[ \frac{4a - 2}{2a^2 -a} = \frac{2(2a - 1)}{a(2a -1)} = \frac{2}{a} \]

OpenStudy (anonymous):

yes that must be it. Thank you very much :)

OpenStudy (anonymous):

Sure np